# Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and Noether-Lefschetz theory of $\mathcal{A}_{g}$

### v1

@@ 5. Calculations
$\begin{split}\langle(2,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=-\frac{t_{1}+t_{2}}{18t_{1}t_{2}}\mathsf{Tr}_{2}^{(2)}-\frac{t_{1}+t_{2}}{16t_{1}t_{2}}\mathsf{Tr}_{3}^{(2,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{4}^{(2,1,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{5}^{(2,1,1,1)}\\ &=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}\,,\end{split}$
$\begin{split}\langle(2,2,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{775t_{1}^{2}+733t_{1}t_{2}+775t_{2}^{2}}{1200(t_{1}+t_{2})}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-10t_{1}^{2}-13t_{1}t_{2}-10t_{2}^{2}}{240t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-3}{200(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}+\frac{-25t_{1}^{2}-68t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}\\[5.0pt] &+\frac{50t_{1}^{2}-139t_{1}t_{2}+50t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{-25t_{1}^{2}-44t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,,\end{split}$
$\begin{split}\langle(3,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{175t_{1}^{2}+77t_{1}t_{2}+175t_{2}^{2}}{300(t_{1}+t_{2})}+\frac{-225t_{1}^{2}-532t_{1}t_{2}-225t_{2}^{2}}{3600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{1}{120(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}\\[5.0pt] &+\frac{-3}{100(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{-25t_{1}^{2}-47t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-2}{75(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}\\[5.0pt] &+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,,\end{split}$
$\begin{split}\langle(3,2)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{865t_{1}^{2}+1556t_{1}t_{2}+865t_{2}^{2}}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-158t_{1}^{2}-91t_{1}t_{2}-158t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{20t_{1}^{2}+79t_{1}t_{2}+20t_{2}^{2}}{450(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}+\frac{-1}{300(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}\\[5.0pt] &+\frac{35t_{1}^{2}-77t_{1}t_{2}+35t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-25t_{1}^{2}-77t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{1}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}+\frac{-25t_{1}^{2}-39t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}\\[5.0pt] &+\frac{-23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}+\frac{-4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,,\end{split}$
$\begin{split}\langle(4,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{-430t_{1}^{2}+743t_{1}t_{2}-430t_{2}^{2}}{2400(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-461t_{1}^{2}+503t_{1}t_{2}-461t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{925t_{1}^{2}+1771t_{1}t_{2}+925t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}\\[5.0pt] &+\frac{-100t_{1}^{2}-223t_{1}t_{2}-100t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}+\frac{-113t_{1}t_{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-1}{50(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{-25t_{1}^{2}-63t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}\\[5.0pt] &+\frac{1}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}+\frac{-11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,,\end{split}$
$\begin{split}\langle(5)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{t_{1}t_{2}(110t_{1}^{2}-1391t_{1}t_{2}+110t_{2}^{2})}{600(t_{1}+t_{2})}+\frac{-75t_{1}^{2}-53t_{1}t_{2}-75t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-50t_{1}^{2}+77t_{1}t_{2}-50t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{75t_{1}^{2}+53t_{1}t_{2}+75t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-5t_{1}^{2}-16t_{1}t_{2}-5t_{2}^{2}}{120t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(5)}\\[5.0pt] &+\frac{175t_{1}^{2}+94t_{1}t_{2}+175t_{2}^{2}}{150(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{44t_{1}^{2}+45t_{1}t_{2}+44t_{2}^{2}}{120(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{-60t_{1}^{2}-53t_{1}t_{2}-60t_{2}^{2}}{200(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,1,1)}\\[5.0pt] &+\frac{15t_{1}^{2}+23t_{1}t_{2}+15t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,.\end{split}$
@@ Appendix A On symmetric functions
$\mathsf{SymDf}\subseteq\mathbb{K}[\mathsf{Ds}][1/\Delta]$

### v2

@@ 5. Calculations
$\langle(2,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,-\frac{t_{1}+t_{2}}{18t_{1}t_{2}}\mathsf{Tr}_{2}^{(2)}-\frac{t_{1}+t_{2}}{16t_{1}t_{2}}\mathsf{Tr}_{3}^{(2,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{4}^{(2,1,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{5}^{(2,1,1,1)}$
$=$ $\,\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}\,$,
$\langle(2,2,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{775t_{1}^{2}+733t_{1}t_{2}+775t_{2}^{2}}{1200(t_{1}+t_{2})}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-10t_{1}^{2}-13t_{1}t_{2}-10t_{2}^{2}}{240t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}$
$+\frac{-3}{200(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}+\frac{-25t_{1}^{2}-68t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}$
$+\frac{50t_{1}^{2}-139t_{1}t_{2}+50t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{-25t_{1}^{2}-44t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,$,
$\langle(3,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{175t_{1}^{2}+77t_{1}t_{2}+175t_{2}^{2}}{300(t_{1}+t_{2})}+\frac{-225t_{1}^{2}-532t_{1}t_{2}-225t_{2}^{2}}{3600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{1}{120(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}$
$+\frac{-3}{100(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{-25t_{1}^{2}-47t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-2}{75(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}$
$+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,$,
$\langle(3,2)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{865t_{1}^{2}+1556t_{1}t_{2}+865t_{2}^{2}}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-158t_{1}^{2}-91t_{1}t_{2}-158t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{20t_{1}^{2}+79t_{1}t_{2}+20t_{2}^{2}}{450(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}+\frac{-1}{300(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}$
$+\frac{35t_{1}^{2}-77t_{1}t_{2}+35t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-25t_{1}^{2}-77t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{1}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}+\frac{-25t_{1}^{2}-39t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}$
$+\frac{-23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}+\frac{-4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,$,
$\langle(4,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{-430t_{1}^{2}+743t_{1}t_{2}-430t_{2}^{2}}{2400(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-461t_{1}^{2}+503t_{1}t_{2}-461t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{925t_{1}^{2}+1771t_{1}t_{2}+925t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}$
$+\frac{-100t_{1}^{2}-223t_{1}t_{2}-100t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}+\frac{-113t_{1}t_{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-1}{50(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{-25t_{1}^{2}-63t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}$
$+\frac{1}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}+\frac{-11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}$
$+\frac{-7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,$,
$\langle(5)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{t_{1}t_{2}(110t_{1}^{2}-1391t_{1}t_{2}+110t_{2}^{2})}{600(t_{1}+t_{2})}+\frac{-75t_{1}^{2}-53t_{1}t_{2}-75t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-50t_{1}^{2}+77t_{1}t_{2}-50t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}$
$+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{75t_{1}^{2}+53t_{1}t_{2}+75t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-5t_{1}^{2}-16t_{1}t_{2}-5t_{2}^{2}}{120t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(5)}$
$+\frac{175t_{1}^{2}+94t_{1}t_{2}+175t_{2}^{2}}{150(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{44t_{1}^{2}+45t_{1}t_{2}+44t_{2}^{2}}{120(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{-60t_{1}^{2}-53t_{1}t_{2}-60t_{2}^{2}}{200(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,1,1)}$
$+\frac{15t_{1}^{2}+23t_{1}t_{2}+15t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,.$
@@ Appendix A On symmetric functions
$\mathsf{SymDf}\subseteq\mathbb{K}[\mathsf{Ds}][1/\Delta]$ .